### v1

# SLMRec: Empowering Small Language Models for Sequential Recommendation

Abstract
The sequential Recommendation (SR) task involves predicting the next item a user is likely to interact with, given their past interactions. The SR models examine the sequence of a user’s actions to discern more complex behavioral patterns and temporal dynamics. Recent research demonstrates the great impact of LLMs on sequential recommendation systems, either viewing sequential recommendation as language modeling or serving as the backbone for user representation. Although these methods deliver outstanding performance, there is scant evidence of the necessity of a large language model and how large the language model is needed, especially in the sequential recommendation scene. Meanwhile, due to the huge size of LLMs, it is inefficient and impractical to apply a LLM-based model in real-world platforms that often need to process billions of traffic logs daily. In this paper, we explore the influence of LLMs’ depth by conducting extensive experiments on large-scale industry datasets. Surprisingly, we discover that most intermediate layers of LLMs are redundant. Motivated by this insight, we empower small language models for SR, namely SLMRec, which adopt a simple yet effective knowledge distillation method. Moreover, SLMRec is orthogonal to other post-training efficiency techniques, such as quantization and pruning, so that they can be leveraged in combination. Comprehensive experimental results illustrate that the proposed SLMRec model attains the best performance using only 13% of the parameters found in LLM-based recommendation models, while simultaneously achieving up to 6.6x and 8.0x speedups in training and inference time costs, respectively.
1 Introduction
Learning temporal interest information is fundamental for sequential recommendation models. Traditional sequential recommendation (TSR) methods [52; 18; 29; 46] focus on the development of intricate sequential encoders, evolving from LSTM and GRU architectures to the self-attention layers and Transformer models. However, the state-of-the-art performance in TSR has hit a plateau, limited by model sizes that usually feature fewer than 0.1 billion parameters.
Recently, Large Language Models (LLMs) [1; 47; 3] have made significant advancements in various aspects by scaling the size of the training data or the model’s architecture. Building upon the scaling laws delineated in prior research [30; 20], it endows Large Language Models (LLMs) with enhanced expressivity, culminating in superior performance benchmarks. Naturally, a burgeoning trend among contemporary LLM-based recommendation architectures has raised concerns. The current LLM-based recommender system can be classified as 1) generation-based approaches, e.g., P5 [12; 54], CoLLM [57] and LLaRa [38]; 2) embedding-based approaches such as E4SRec [35], CLLM4Rec [59] and Lite-LLM4Rec [50]. As shown in Fig. 1, generation-based approaches (G-LLMRec) encode an item as a token and formulate the sequential recommendation as the next token prediction task. By contrast, embedding-based approaches (E-LLMRec) regard the last hidden representation as user representation and learn an external adapter to compute user-item preference. The adoption of LLMs has vastly driven the development of sequence recommendation tasks, bringing an improvement of nearly 20% against the TSR model on the benchmark [35; 38; 50]. This arouses the following research motivation for this work.
$\bullet$ So far on the SR scene, little literature studies how large the language Model is needed. Recent studies [37; 14; 40] on the NLP domain suggest a high degree of redundancy in the LLM’s model architecture. Since the ID information has not been explicitly learned during the LLM’s training process, we also want to find out whether the scaling law of LLMs is also applicable to the SR task, specifically, how the relation is between the size of LLMs and their performance on SR task.
$\bullet$ Despite the large performance gain, the LLM-based SR solutions also escalate the model size significantly, e.g., nearly 70 times greater parameters compared with TSR models (from 0.1B to 7B+). Even within the parameter-efficient training technique [24], the paradigm still poses a significant challenge for real-world sequential recommendation use cases, where billions of traffic logs every day and potential new items need to be processed constantly. This disparity imposes strict hardware demands and makes it both inefficient and infeasible to deploy an LLM-based SR model.
Figure 1: This overview compares traditional sequential recommendation (TSR) methods with LLM-based recommendation (LLMRec) methods. Here, $h_{u}$ and $h_{i}$ represent the user and item representations, respectively. In contrast to G-LLMRec methods [12; 54; 57; 38], E-LLMRec approaches [35; 59; 50] adhere to the TSR prediction framework. These methods leverage LLMs as feature extractors in the manner of BERT [8], diverging from the generative focus of G-LLMRec.
Our contributions. This paper presents our initial attempt to reassess the need for large language models in sequential recommendation. To explore the reasons for the significant improvement, we conduct a series of experiments on large-scale industry datasets to investigate the effects of reducing the number of parameters during the training and inference stages on overall performance. From the empirical results, we found some profound insights that the improvement of the rise of the model parameters is not consistent. Meanwhile, it reveals that some layers of LLMs are redundant in the downstream recommendation task, similar to some recent findings in NLP domains [40; 14].
Motivated by these findings, we empower small language models for the sequential recommendation, named SLMRec. We adopt the vanilla knowledge distillation approach [13; 6; 2] to align the representation knowledge. Moreover, multiple supervision signals are crafted to steer the student model toward acquiring task-aware knowledge within its hidden representations. Additionally, SLMRec operates without the need for any supplementary model design elements and is compatible with other quantization and pruning techniques utilized within Large Language Models (LLMs).
Extensive experiments have revealed that SLMRec, with a model size of less than 1 billion parameters, can deliver performance that is remarkably competitive with baselines using LLMs sized over 7 billion parameters. Furthermore, SLMRec achieves up to 6.6x/8.0x speedup in terms of training/inference time costs against LLM-based recommendation models. Furthermore, we present the results of SLMRec employing online knowledge distillation, showing its competitive performance as well.
[…]
2 Motivation
As described above, here we try to explore the effectiveness of LLMs in recommendation via decreasing the parameters of popular LLMs (i.e., LLaMa-7B) and observe the change in performance.
Evaluation Protocol. In the motivational experiment, we select SASRec as a traditional sequential recommendation baseline due to its performance [32]. We adopt [^1] To accelerate and follow the prediction head of traditional SR methods, we remove the original softmax layer and apply the dot product of user representation and item representation to obtain user-item prediction score. embedding-based method [35] as the baseline, named E4SRec, to easily generate the ranking for the full/sampled list of items. As shown in Fig. 2, a pre-trained embedding layer learned from SASRec is used to obtain the sequential item embedding. Then we concatenate the item embeddings with the prompt embeddings obtained after the tokenization. After encoding of stacked attention blocks of LLM, we regard the representation of the last layers as the user representation. Then, we follow the TSR methods to calculate the inner product of user embeddings and item embeddings from the pre-trained embedding layer to serve as the score for the user-item pair. Also, cross-entropy loss and fully candidate item are utilized for the optimization to achieve best results [53; 43]. To reduce both computational demands and processing time, LoRA [24] is used to update a comparatively smaller set of parameters. Besides, to generate an unbiased evaluation for fair comparison [33; 58], we randomly sampled 999 negative items, which were items not interacted with by the user, along with 1 positive item that served as the ground-truth interaction. To obtain large-scale industry data, we use the Amazon 18 version222 https://nijianmo.github.io/amazon/index.html dataset in this paper. More details are shown in Section 5.
[…]
3 Preliminaries
Figure 3: The overview of $\mathrm{SLMRec}$. A layer-wise knowledge distillation approach is applied to align the representation knowledge by grouping the layer into serveral blocks. The teacher and student model share a similar E-LLMRec model architecture. Multiple supervision signals are introduced to steer the student model toward acquiring fine-grained task-aware knowledge.
In this study, rather than constructing complex additional structures, we slightly modify existing E-LLMRec methods for our purposes. Initially, we delineate the E-LLMRec model that we employ for sequential recommendation tasks.
Model structure. The E-LLMRec models capitalize on an ID embedding layer from TSR models such as BERT4Rec, SASRec, and GRU4Rec, which is pre-trained on a designated dataset [46; 29; 18]. The objective of sequential recommendation is to forecast subsequent items utilizing the user action sequence $S=(i_{1},i_{2},...,i_{T})$, a sequence that is either truncated or padded to maintain uniform length. Through truncation and padding, we derive the user’s action sequence mask, serving as the attention mask in LLMs (Large Language Models). The fixed-length sequence $S\in\mathbb{R}^{T}$ is translated into a sequential representation $\mathbf{S}\in\mathbb{R}^{T\times d_{0}}$ via the pre-trained ID embedding layer. A linear transformation is then applied to upscale the representation from a lower dimension $d_{0}$ to a higher dimension $d_{1}$ suitable for the hidden layers within the LLMs.
Upon defining the prompt template, the tokenization layer within the LLMs processes the natural language input into corresponding text embeddings and their associated attention masks. These embeddings and attention masks, derived from both the ID sequence and the text, are then introduced into the LLM decoder. The final temporal output $\mathbf{h}_{M}$ from the last layer of the decoder is inferred as the user representation and subsequently mapped through a linear layer to condense the dimensionality from $d_{1}$ back to $d_{0}$. Finally, user-item interaction predictions $\bar{p}$ are inferred by executing a dot product between the user and item representations. The learning process is refined through the application of a cross-entropy loss.
$\hat{p}_{i}=\frac{e^{\bar{p_{i}}}}{\sum\limits_{j\in I}e^{\bar{p_{j}}}};\;\;\;\mathcal{L}_{ce}=-\sum\limits_{u\in U,i\in I}y_{ui}log(\hat{p}_{i}).$ (1)
where $U$ and $I$ denote the whole user set and item set. $y_{ui}$ denotes user-item interaction label.
Knowledge Distillation. Knowledge distillation is a technique aimed at transferring knowledge from a sophisticated teacher model to a more streamlined student model [19]. We represent the teacher by $f_{t}(\Theta^{t})$ and the student by $f_{s}(\Theta^{s})$. We aim to solve the following optimization problem:
$\min_{\Theta^{s}}[\mathcal{L}_{ce}(\Theta^{s})+\mathcal{D}_{kd}(\Theta^{t},\Theta^{s})].$ (2)
Here, $\mathcal{D}_{kd}(\Theta^{t},\Theta^{s})$ signifies the knowledge distillation loss, which quantifies the discrepancies between the teacher and the student models. A prevalent method involves employing the KL divergence to evaluate the divergence between the logits produced by both models. One well-established training schema is known as offline distillation, wherein the teacher is fully trained beforehand and remains unchanged, while the student is refined based on the criteria outlined in Eq. 6. With a designated training set, the teacher model $\Theta^{t}$ is initially trained by minimizing the cross-entropy loss $\mathcal{L}_{ce}$.
4 SLMRec
In this work, we do not conduct logits-based knowledge distillation since we want the student model to have the same ability to encode hidden representations rather than making similar predictions as the teacher. Therefore, we conduct feature distillation among every few layers. Considering the teach model has $M$ stacked decoder layers and student model has $N$ stacked decoder layers, we design several feature regularizars to conduct knowledge distillation between hidden representation from teach model $\mathbf{H}^{t}=\{\mathbf{h}^{t}_{m},...,\mathbf{h}^{t}_{m*i},...,\mathbf{h}^{t}_{b}\}$ and student model $\mathbf{H}^{s}=\{\mathbf{h}^{s}_{n},...,\mathbf{h}^{s}_{n*i},...,\mathbf{h}^{s}_{b}\}$ each few layers. Each $m$ layer of the teacher model and each $n$ layer of the student model are regarded as a block. The block number is $b$. In this work, we consider deeper LLaMa as the teacher model, and shallow LLaMa as the student model, which has the same hidden dimension between $\mathbf{H}^{t}\in\mathbb{R}^{b\times d_{1}}$ and $\mathbf{H}^{s}\in\mathbb{R}^{b\times d_{1}}$.
Feature Similarity. To regulate the alignment of feature directions between the teacher and student models, we employ a cosine similarity-based loss term. Formally, it is described by the equation:
$\mathcal{D}_{cos}(\Theta^{t},\Theta^{s})=\frac{1}{b}\sum\limits_{k\in b}\frac{\mathbf{h}^{t}_{k}\cdot\mathbf{h}^{s}_{k}}{\|\mathbf{h}^{t}_{k}\|_{2}\cdot\|\mathbf{h}^{s}_{k}\|_{2}}.$ (3)
Feature Norm Regularization. In addition, we introduce a straightforward regularization term designed to minimize the L2 distance between the hidden representations of the teacher and student models. It is mathematically formulated as:
$\mathcal{D}_{norm}(\Theta^{t},\Theta^{s})=\frac{1}{b}\sum\limits_{k\in b}\|\mathbf{h}^{t}_{k}-\mathbf{h}^{s}_{k}\|_{2}^{2}.$ (4)
Multiple Supervision. Furthermore, we employ multiple supervision strategies to steer the student model toward assimilating specific aspects of recommendation-related knowledge. For each representation, we learn additional adapters ( $W_{t}$ ) to reduce the dimension. The modified prediction ( $\hat{p}^{mp}$ ) can be acquired as described by Eq. 1:
$\mathcal{L}_{ms}(\Theta^{s},W_{t})=\frac{1}{b}\sum\limits_{k\in b}\mathcal{L}_{ce}(y,\hat{p}^{mp}).$ (5)
Total Loss. Integrating the aforementioned distillation losses, the composite objective function for training the student model is given by:
$\min_{\Theta^{s},W_{t}}[\mathcal{L}_{ce}(\Theta^{s})+\lambda_{1}\mathcal{D}_{cos}(\Theta^{t},\Theta^{s})+\lambda_{2}\mathcal{D}_{norm}(\Theta^{t},\Theta^{s})+\lambda_{3}\mathcal{L}_{ms}(\Theta^{s},W_{t})].$ (6)
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are hyperparameters that govern the contribution of each term to the knowledge distillation process.
5 Experiments
In this section, we present extensive experiments to demonstrate the effectiveness of $\mathrm{SLMRec}$, aiming to answer the following three research questions (RQs).
$\bullet$ RQ1: How does the performance of our proposed SLMRec model compare to LLM-based recommendation models when evaluated on a large-scale industry dataset?
$\bullet$ RQ2: What is the comparative efficiency and runtime of our SLMRec model against the G-LLMRec and E-LLMRec models?
$\bullet$ RQ3: Whether the proposed three knowledge regularizers work?
$\bullet$ RQ4: Is it feasible to train our SLMRec model without utilizing a pre-trained teacher model?
5.1 Experiment Setup
Table 1: Statistics on the Amazon datasets.
Dataset $\left|\mathcal{U}\right|$ $\left|\mathcal{V}\right|$ $\left|\mathcal{E}\right|$ Density
Cloth 1,219,678 376,858 11,285,464 0.002%
Movie 297,529 60,175 3,410,019 0.019%
Music 112,395 73,713 1,443,755 0.017%
Sport 332,447 12,314 146,639 0.008%
  • $\left|\mathcal{U}\right|$, $\left|\mathcal{V}\right|$, $\left|\mathcal{E}\right|$ denote the number of user, item and ratings, respectively.
For our experimental evaluation, we utilize data from the clothing, movies, music, and sports categories within the extensive, industry-scale Amazon18 dataset  [^3] https://nijianmo.github.io/amazon/index.html. Statistics of the datasets are shown in Table 1. In all datasets, we interpret any rating above 3 as implicit feedback, indicating user interaction with the item, and employ timestamps to establish the chronological order of actions. We eliminate users and items that have fewer than 5 associated actions to ensure sufficient data density. The historical sequence of interactions for each user is divided into three segments: (1) the most recent interaction is reserved for testing, (2) the second most recent for validation, and (3) all preceding interactions are used for training. Based on the ranking results, we utilize the typical $top$-$N$ metrics hit rate (HR@{1, 5, 10}), normalized discounted cumulative gain (NDCG@{5,10}) [26] and Mean Reciprocal Rank (MRR) [45] to evaluate the model performance. For all the metrics, higher values indicate better performance. Models that achieve the highest MRR performance on the validation set, including ours and other baseline models, will be preserved for subsequent performance evaluation on the test set. In order to ensure an unbiased evaluation, we adopt the methodology employed in previous works [33; 58], wherein we randomly select 999 negative items (i.e., items that the user has not interacted with) and combine them with 1 positive item (i.e., a ground-truth interaction) to form our recommendation candidates for the ranking test. Detailed hyperparameters of our model in each dataset can be obtained in Appendix A.1.
5.2 Performance Comparisons
Compared Methods. We compare our method with three classes of baselines: (1) Single-domain sequential recommendation methods, i.e., BERT4Rec [46], GRU4Rec [18] and SASRec [29]. (2) G-LLMRec method: Open-P5LLaMa [^4] For Open-P5, we adopt the version of LLaMa as the foundation model in their code repository implementation to ensure the best results are achieved. [54]. (3) E-LLMRec method: E4SRec* [35]. A detailed introduction to these baselines can be found in Appendix A.2. It should be noted that we did not select various G-LLMRec methods or E-LLMRec methods as baselines. This is because the differences between each LLM-based method are minimal, and our model is a universal approach that is not confined to a specific model type. Our primary focus is to improve the efficiency of language model utilization. Hence, we opted to select one generative-based method (Open-P5) and one embedding-based method (E4SRec) as our baselines.
Table 2: Experimental results (%) on the Cloth and Movie dataset. The missing MRR value of Open-P5 is unavailable due to the time complexity constrictions. The number on the left of the arrow is the layers $N$ of the student model. The left number on the right of the arrow is the layers $M$ of the teacher model.
Methods Cloth Dataset Movie Dataset Average
Improv.
HR NDCG MRR HR NDCG MRR
@1 @5 @10 @5 @10 @1 @5 @10 @5 @10
GRU4Rec [18] 13.79 15.46 16.83 14.64 15.08 15.15 10.56 19.47 25.21 15.11 16.96 15.46 -20.26
BERT4Rec [46] 13.60 14.66 15.55 14.14 14.43 14.59 9.68 14.91 17.98 12.40 13.38 12.74 -29.84
SASRec [29] 13.08 16.94 20.26 15.01 16.08 15.76 5.57 16.80 26.85 11.17 14.42 12.08 -25.19
Open-P5LLaMa [54] 14.13 17.68 19.74 17.02 16.40 - 12.66 21.98 27.24 17.13 19.81 - -9.96
E4SRec* [35] 16.71 19.45 21.86 18.09 18.86 18.77 14.74 23.79 29.09 19.45 21.16 19.74 0.00
E4SRec8[35] 15.30 18.54 21.29 16.91 17.79 17.60 13.32 22.49 28.57 17.99 19.94 18.46 -5.90
E4SRec4[35] 14.58 18.05 20.92 16.32 17.25 17.01 11.80 21.54 28.02 16.73 18.82 17.20 -10.24
$\mathrm{SLMRec}_{8\leftarrow 32}$ 16.56 19.05 21.33 17.79 18.53 18.48 15.18 23.93 29.30 19.69 21.41 20.06 -0.17
$\mathrm{SLMRec}_{4\leftarrow 32}$ 14.86 18.03 20.70 16.45 17.30 17.12 13.70 22.73 28.44 18.37 20.21 18.74 -6.56
$\mathrm{SLMRec}_{4\leftarrow 8}$ 16.10 18.85 21.33 17.48 18.28 18.17 14.83 23.08 28.02 19.08 20.67 19.45 -2.55
+$\mathcal{D}_{norm}$:
$\mathrm{SLMRec}_{4\leftarrow 8}$ 16.28 19.12 21.75 17.69 18.53 18.40 14.86 23.89 30.22 19.36 21.39 19.84 -0.37
+$\mathcal{L}_{ms}$:
$\mathrm{SLMRec}_{4\leftarrow 8}$ 16.85 19.05 20.93 17.96 18.57 18.59 15.05 23.48 28.60 19.40 21.05 19.76 -0.85
+$\mathcal{D}_{norm}$+$\mathcal{L}_{ms}$:
$\mathrm{SLMRec}_{4\leftarrow 8}$ 16.69 19.47 21.90 18.07 18.85 18.74 15.29 24.25 30.19 19.90 21.82 20.36 +1.49
Table 3: Experimental results (%) on the Music and Sport dataset.
Methods Music Dataset Sport Dataset Average
Improv.
HR NDCG MRR HR NDCG MRR
@1 @5 @10 @5 @10 @1 @5 @10 @5 @10
GRU4Rec [18] 1.89 3.22 4.36 2.57 2.93 3.08 5.26 7.75 9.64 6.52 7.13 7.08 -41.93
BERT4Rec [46] 2.10 3.16 4.04 2.64 2.92 3.11 4.81 6.70 7.98 5.79 6.20 6.26 -46.82
SASRec [29] 1.82 5.72 8.64 3.79 4.72 4.51 4.70 8.43 11.21 6.59 7.48 7.24 -30.60
Open-P5LLaMa [54] 4.35 8.12 10.18 6.74 7.83 - 5.49 8.5 11.28 6.92 7.65 - -12.62
E4SRec* [35] 5.62 9.29 11.64 7.50 8.26 7.98 6.40 9.67 12.35 8.05 8.91 8.70 0.00
E4SRec8[35] 5.46 8.86 11.14 7.21 7.95 7.74 5.48 8.63 11.34 7.06 7.93 7.76 -7.48
E4SRec4[35] 5.33 8.75 10.94 7.08 7.78 7.59 5.41 8.65 11.35 7.04 7.91 7.72 -8.54
$\mathrm{SLMRec}_{8\leftarrow 32}$ 5.65 8.89 11.34 7.30 8.09 7.90 5.43 8.77 11.45 7.11 7.97 7.78 -6.39
$\mathrm{SLMRec}_{4\leftarrow 32}$ 5.26 8.55 10.88 6.96 7.71 7.50 5.76 8.91 11.45 7.35 8.17 8.02 -7.56
$\mathrm{SLMRec}_{4\leftarrow 8}$ 5.62 8.78 11.09 7.23 7.97 7.81 6.25 9.25 11.68 7.76 8.54 8.41 -3.55
+$\mathcal{D}_{norm}$:
$\mathrm{SLMRec}_{4\leftarrow 8}$ 5.95 9.26 11.68 7.65 8.42 8.23 6.61 9.82 12.34 8.24 9.05 8.87 +1.97
+$\mathcal{L}_{ms}$:
$\mathrm{SLMRec}_{4\leftarrow 8}$ 5.69 8.94 11.18 7.36 8.08 7.91 6.51 9.39 11.80 7.96 8.74 8.62 -1.75
+$\mathcal{D}_{norm}$+$\mathcal{L}_{ms}$:
$\mathrm{SLMRec}_{4\leftarrow 8}$ 5.72 9.15 11.45 7.48 8.21 8.03 6.62 9.83 12.34 8.25 9.06 8.89 +0.81
Quantitative Results (RQ1). Tables 23 showcase the quantitative comparison of four large-scale sequential recommendation datasets. From our analysis, we have several insightful observations: (1) LLM-based recommendation methods exhibit substantial improvements over traditional sequential recommendation (TSR) methods, primarily due to their enhanced modeling capacity which adeptly extracts informative sequential interest patterns. (2) Our model, $\mathrm{SLMRec}_{4\leftarrow 8}$, outperforms the teacher model E4SRec8* by leveraging knowledge distillation within the hidden layers. By refraining from applying this constraint prior to the prediction phase, we enable the final representation to organically gravitate towards the label—yielding an approximate 8% enhancement in performance in comparison to the teacher model. (3) Introducing vanilla knowledge distillation techniques into LLMRec, without altering the model structure, allows $\mathrm{SLMRec}_{4\leftarrow 8}$ to achieve a marginally superior performance compared to E4SRec32*. This suggests that small language models equipped with efficacious training strategies can rival, or even exceed, larger language models in the sequential recommendation task.
Model Efficiency (RQ2). We report the time efficiency and parameters of comparative baselines and our model in Table 4. All time and parameter metrics represent the average across the four datasets reported. Inference time evaluates the prediction ranking among 1,000 candidate items for each user. Detailed training and inference times for each dataset are provided in Appendix A.3. The Open-P5, an LLMRec model based on generative methods, offers a reasonable training duration. Yet, during the inference phase, it becomes considerably time-consuming (4942 hours) as it necessitates generating a substantial pool of candidate items (for instance, 1000). Owing to the intrinsic workings of generative LLMs, employing generation-based LLMRec models for the comprehensive ranking of extensive item sets is not advised. Our model outperforms E4SRec with enhanced efficiency, maintaining only 13% and 14% in E4SRec’s parameters for training and inference, respectively. Moreover, our SLMRec demonstrates a remarkable gain in speed, being 6.6 times faster during training and 8.0 times quicker in inference than E4SRec.
Ablation Study (RQ3). In all experimental results, SLMRec, when enhanced with various knowledge regularizers (namely $\mathcal{D}_{cos},\mathcal{D}_{norm}$ and $\mathcal{L}_{ms}$), demonstrates improved performance. The regularizers $\mathcal{D}_{cos}$ and $\mathcal{D}_{norm}$ aid SLMRec in aligning its intermediate representations with those of the teacher model, thereby endowing it with more potent representational extraction capabilities. Meanwhile, $\mathcal{L}_{ms}$ steers the model to assimilate domain knowledge pertinent to recommendation systems within its preliminary layers.
Table 4: Efficiency comparison of Open-P5, E4SRec, and our SLMRec in terms of epoch-wise training time (hours), inference time (hours), number of training parameters (B) and inference parameters (B). These comparisons were conducted on a machine with an A100 GPU. The training batch size for all models was standardized at 256. During inference, E4SRec and SLMRec utilized a batch size of 512, whereas Open-P5’s inference was performed with a batch size of 1.
Method Tr time(h) Inf time(h) Tr params (B) Inf params (B)
Open-P5LLaMa 0.92 4942 0.023 7.237
E4SRec* 3.95 0.415 0.023 6.631
SLMRec4←8 0.60 0.052 0.003 0.944
5.3 Model Study
Study of Online KD (RQ4). In our methodology, we first train the teacher model on downstream recommendation tasks and then train the student model through knowledge distillation, which is an offline knowledge distillation technology. In this section, we demonstrate that we can train both the teacher model and SLMRec together on downstream recommendation tasks, which constitutes an online knowledge distillation. Under this setting, we are able to achieve comparative results.
Study of block number $b$. We also conducted experiments to investigate the effect of block number $b$. As shown in Figure  4, when $b$ is set to 4, our model achieves the best performance. When $b$ is set to 1 or 2, the feature constraint imitation for each block within SLMRec is diminished relative to the teacher model, resulting in a decline in performance.
(a) Online KD
(b) Online KD
(c) Block Number
(d) Block Number
Figure 4: Model study of online KD and block number $b$ in the music dataset.
6 Related Work
In this section, we introduce the most related background and scientific investigations to this work, which are roughly divided into five categories, i.e., 1) Sequential Recommendation, 2) Knowledge Distillation (KD) 3) Depth-wise Knowledge of LLMs, 4) Model Pruning, and 5) Parameter-Efficient Fine-Tuning (PEFT).
Sequential Recommendation. Traditional Sequential Recommendation (TSR) methods [52; 18; 46; 29] primarily focus on developing various temporal encoders to capture short- and long-term user interests. Following the triumph of large language models (LLMs), researchers have begun leveraging open-source LLMs [47] to construct their recommendation systems [27; 5; 51]. G-LLMRec methods [12; 54; 57; 38] generate the next item based on historical sequences, while E-LLMRec approaches [35; 59; 50] use LLMs as feature extractors to learn user representations for prediction. More recently,  [55] introduces a generative sequential framework scalable up to GPT-3 dimensions, trained from scratch with 100 billion examples and utilizing 256 H100 GPUs. LLM-based recommendation systems frequently outperform TSR models by a margin of 20% [35; 38; 50], also increasing the parameters by nearly 100 times compared to TSR models. Therefore, the deployment of LLMRec models in real-world platforms is heavily constrained by computational resources.
Knowledge Distillation (KD). Training a smaller “student" model on the distribution predicted by a large “teacher" model is known as a powerful knowledge distillation technique [19]. The fundamental insight behind this is to transform the knowledge and capabilities of the teacher into more compact, compressed, and possibly skill-specific representations [28; 15]. For those cases when the student only has access to the output tokens generated by the teacher, another way of KD is data distillation [9; 36; 11; 23]. This technique first generates high-quality synthetic data by prompting the larger teacher model. The synthetic data are then used to enhance the student’s capabilities by fine-tuning. Our work lies in the former series, which performs a simple yet effective layer-skipping approach to transfer those useful representations of the teacher to the student.
Depth-wise Knowledge of LLMs. The recent community interest stems from how linguistic properties and knowledge are encoded in language models.  [41; 7] emphasize that knowledge localizes within the middle or final layers. On the other hand,  [16] attempts to perform knowledge editing and concludes that information may be stored non-locally across layers. What’s more,  [40; 14] share a similar view that current pretraining methods are not properly leveraging the parameters in the deeper layers of the network or that the shallow layers play a critical role in storing knowledge. By contrast, we are the first to investigate which part of knowledge on the LLMs plays a key role, especially in the sequential recommendation scene.
Model Pruning. Model Pruning is a fundamental approach for reducing the size of a well-trained large model by removing unimportant parameters [17]. Recent work has focused on applying pruning methods to the Transformer architecture [48]. These works have studied different components of the model architecture for pruning, including dropping attention heads [49; 42], dropping layers  [10; 56; 31; 44], dropping hidden states [21], replacing sparse weight matrices with smaller dense ones [4], and combinations of these solutions. By contrast, our work performs layer removal through simple knowledge distillation, rather than more complex pruning techniques.
Parameter-Efficient Fine-Tuning (PEFT). PEFT emerges as a novel technique for tailoring Large Language Models (LLMs) to specific tasks while ensuring minimal computational and memory costs [22; 34; 25; 39]. In this work, we combine our method with the Low-Rank Adapters (LoRA) [25] to reduce the memory and computation of the knowledge distillation process. Specifically, we freeze the pre-trained model and only tune a small set of additional trainable parameters.
7 Conclusions and Outlooks
This paper explores the effectiveness of large language models in sequential recommendation. Motivated by empirical insights, we adopt vanilla knowledge distillation methods to improve the performance of small language models. With only 13% parameters and yields 8x acceleration against the LLMRec baseline, our SLMRec achieves slightly better performance. On top of our technical contributions, we believe the results in this paper could shed light on a new promising direction for building effective and efficient recommenders based on LLMs, which is largely under-explored.
Future Work. This work concentrates on enhancing the efficiency of Large Language Model (LLM) utilization in the sequential recommendation. A notable limitation is the model’s inability to adapt to new scenarios through few-shot learning. When confronted with a fresh dataset or new traffic logs from the platform, the model requires retraining from the entire dataset. In contrast, LLMs have demonstrated promising results in adapting to downstream language tasks using few-shot learning approaches. Looking ahead, we intend to investigate the incorporation of incremental learning into LLM-based recommendations to bolster the model’s transferability. Additionally, integrating auxiliary linguistic and visual information of users and items into the LLMRec model may offer further improvements in its adaptability to new scenarios.
Appendix A Appendix
A.1 Training Details
In Table 5, we provide hyper-parameters in our training stage. Our implementation is based on Huggingface Transformers  [^5] https://github.com/huggingface/transformers. The input and intermediate hidden dimension in the feed-forward network is 4096. We use mixed precision training and train on 1*80G Nvidia A100 GPU.
Table 5: Hyper-parameter (HP) settings of our method on each dataset.
HP0 Cloth Movie Music Sport
adam_beta1 0.9 0.9 0.9 0.9
adam_beta2 0.999 0.999 0.999 9.999
adam_epsilon 1e-8 1e-8 1e-8 1e-8
learning_rate 0.003 0.001 0.002 0.002
logging_steps 1 1 1 1
lr_scheduler_type cosine cosine cosine cosine
max_grad_norm 1.0 1.0 1.0 1.0
max_steps 1500 -1 800 2000
optimizer 0 adamw_torch0 0 adamw_torch0 0 adamw_torch0 0 adamw_torch0
save_strategy steps steps steps steps
save_steps 50 100 100 100
eval_steps 50 100 100 100
warmup_steps 50 50 100 50
$\lambda_{1}$ 1.0 1.0 1.0 1.0
$\lambda_{2}$ 0.1 0.1 0.1 0.1
$\lambda_{3}$ 1.0 1.0 0.01 0.1
$b$ 4 4 4 4
A.2 Compared methods
Tranadtional sequential recommendation methods:
GRU4Rec [18] tackles the issue of modeling sparse sequential data while also adapting RNN models to the recommender system. To achieve this, the authors propose a new ranking loss function that is specifically designed for training these models. The implementation of GRU4Rec in PyTorch can be found at the URL  [^6] https://github.com/hungpthanh/GRU4REC-pytorch.
BERT4Rec [46] designs a bidirectional self-attention network to model user behavior sequences. To prevent information leakage and optimize the training of the bidirectional model, a Cloze objective is used to predict the randomly masked items in the sequence by considering both their left and right context. The implementation of BERT4Rec in PyTorch can be found at the URL  [^7] https://github.com/jaywonchung/BERT4Rec-VAE-Pytorch.
SASRec [29] is a self-attention based sequential model that addresses the challenge of balancing model parsimony and complexity in recommendation systems. By using an attention mechanism, SASRec identifies relevant items in a user’s action history and predicts the next item based on relatively few actions, while also capturing long-term semantics like an RNN. This enables SASRec to perform well in both extremely sparse and denser datasets. The implementation of SASRec in PyTorch can be found at the URL  [^8] https://github.com/pmixer/SASRec.pytorch.
LLM-based recommendation methods:
Open-P5 [54] is an open-source platform introduced to catalyze research in LLM-based generative recommender systems. It supports key model architectures like T5 and Llama-2 across diverse public datasets, focusing on sequential and straightforward recommendation tasks. The platform emphasizes the role of item IDs through various indexing methods and offers a customizable, efficient, and standardized environment for developing and assessing recommender systems. The implementation of Open-P5 in PyTorch can be found at the URL   [^9] https://github.com/agiresearch/OpenP5.
E4SRec [35] integrate of Large Language Models (LLMs) into sequential recommendation systems, offering a significant leap in handling item IDs and personalization. In the original paper, they use Softmax layer to output each user-item prediction score. The implementation of E4SRec in PyTorch can be found at the URL  [^10] https://github.com/HestiaSky/E4SRec.
A.3 Model Efficiency
We show the running time of Open-P5, E4SRec, and our SLMRec in each dataset. These comparisons were conducted on a machine with an A100 GPU. The training batch size for all models was standardized at 256. During inference, E4SRec and SLMRec utilized a batch size of 512, whereas Open-P5’s inference was performed with a batch size of 1.
Table 6: Detailed efficiency comparison of Open-P5, E4SRec, and our SLMRec, in terms of training and inference time, on each dataset.
0Method 0 Cloth0 Movie 0 Music0 Sport
0 Tr time (h) 0 Inf time (h) 0 Tr time (h) 0 Inf time (h) 0 Tr time (h) 0 Inf time (h) 0 Tr time (h) 0 Inf time (h)
Open-P5LLaMa 1.36 3554.43 0.36 3504 0.35 3692 1.60 9017
E4SRec* 5.27 0.578 1.90 0.208 1.88 0.216 6.75 0.660
SLMRec4←8 0.97 0.070 0.15 0.030 0.30 0.030 0.98 0.078